A backyard swimming pool holds a volume of 190 cubic yards (yd3) of water. How many of the following dimensions are equivalent to 190 yd3 ? Assume the density of water is 1.00 g/mL


1
Expert's answer
2021-09-18T23:58:19-0400

1yd3= 0.9144m3

190 × 0.9144= 173.74ml

1cm3 = 1ml

173.74 × 106cm3

1000cm3 = 1kg

= 173.74 × 106/103

= 173.74 × 103kg

1kg = 2.2lb

= 173.74 × 103 × 2

= 3.47472 × 105lb